OpenStudy (anonymous):

what is 3/14 times 4/7??

OpenStudy (anonymous):

24/18 you can simplify it.

OpenStudy (anonymous):

It's 1 1/3 simplified

OpenStudy (anonymous):

thanks can u help me o n one more

OpenStudy (anonymous):

what is 13/15 + -1/5

OpenStudy (anonymous):

1. (3*4)/(14*7) 12/98=2/13 2. 13/15+ -1/5 13/15-3/15 10/15=2/3

OpenStudy (anonymous):

thsnks

OpenStudy (anonymous):

one more???

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

It's 2/5

OpenStudy (anonymous):

no 10/5=2 15/5=3 2/3 @Nathan28

OpenStudy (anonymous):

2 3/4 - 1/2

OpenStudy (anonymous):

2*4+3=11 11/4-1/2 11/4-2/4 9/4=2 1/4

OpenStudy (anonymous):

hey private message me

OpenStudy (anonymous):

ok